OpenStudy (anonymous):

PLEASE PLEASE PLEASE HELP ME :( i have an exam. :( A woman bought a bed priced at $22 950. she decided to pay by installment for 18 months using her credit card. the finance charge for the 18 month installment is 1.5% per month. a one-time 5% interest on the list price will be added for the use of a credit card. she could only set aside $1200 each month for the next 18 months. how much downpayment should she make if this is her limitation?

OpenStudy (anonymous):

what i TRIED to do was... 22950(0.05)+22950 so the new price is now 24097.5 then i computed for the finance charge which is 24097.5(0.015)(18) i didn't do 18/12 or convert the 18 months to years since the rate is stated per month... did i do that right?? anyway so i got the finance charge as 6506.325 then i did 1200=(24097.5+6506.325-downpayment)/18 solved for downpayment and i got-- 9003.825 did i do it right?? please someone confirm thank you!!

OpenStudy (anonymous):

help please? anyone?? thank you i'd reaaaally appreciate it!!! @campbell_st @cwrw238 @ganeshie8 @Agent_Sniffles @AccessDenied @him1618 @Preetha @Jonask @nitz @kelly226 @nubeer @nincompoop @Nurali @ByteMe @LonelyandForgotten @BrennaM etcetcetc???

OpenStudy (anonymous):

we can use the formular\[F=x \frac{\left[ (1+i)^n-1 \right]}{i}\] x=1200 i=0.015 n=18 i am not sure about the downpayment what it means

OpenStudy (anonymous):

can you explain the formula you used? anyway downpayment is what you pay in advance so you subtract that from the principal

OpenStudy (anonymous):

this is just the sum formular ,to calculate the sum of all the monthly payments i guess since you never so it probably you cant use it \[S_n=\frac{ a(r^n-1) }{ r-1 }\]

OpenStudy (anonymous):

Do you still need help?

OpenStudy (anonymous):

@Skaematik, yes please!!

OpenStudy (anonymous):

@hartnn @TuringTest ? anyone?

OpenStudy (anonymous):

@Zarkon @cwrw238 @asnaseer @Yahoo! anyone please? :(

OpenStudy (anonymous):

The loan amount including one-time credit charge: $22 950.( 1 + .05 ) = $ 24,097.5 The amount she's able to pay installment in 18 months: PV = $1,200 [1 - 1.015^(-18 ) ] / .015 = $18, 807. 07 Thus down payment: $ 24,097.5 - $18, 807. 07 = ....

OpenStudy (anonymous):

@Chlorophyll can you please explain your 2nd equation/formula? :D

OpenStudy (anonymous):

Do you know the Present Value formula?

OpenStudy (anonymous):

nope we weren't taught that, it's not included for our exam... O:

OpenStudy (anonymous):

but are you really sure we're supposed to use that formula? is there another way without using it? maybe the formula is derived from somewhere? can you tell me where the 1.015 part in the equation came from?

OpenStudy (anonymous):

( 1 + .015 ) = 1.015

OpenStudy (phi):

what formulas have you been taught?

OpenStudy (anonymous):

@phi apr formula, simple and compound interest...?

OpenStudy (phi):

Did you learn the amortization formula ? See http://en.wikipedia.org/wiki/Annual_percentage_rate#Rate_format \[ \text{payment }= \text{Principal }\cdot \frac{i\cdot (1+i)^n}{(1+i)^n-1} \] you could use this formula using i= 0.015, n=18, payment =1200, and solve for the Principal. You get 18807.07 for the principal. Because the amount of the purchase is 24097.5, they need to make a down payment of 24097.50- 18807.07 = 5290.43

OpenStudy (anonymous):

@phi Where does the 5% flat interest due to credit card come in?

OpenStudy (anonymous):

@phi nope not yet :( maybe this isn't included in our exam after all since we don't know that yet. i just got this question from an old test.

OpenStudy (phi):

@Skaematik a one-time 5% interest on the list price will be added for the use of a credit card. This means that you add 5% to the cost of the item, and that is how much you have to pay. It is the opposite of a 5% off sale, where the price is lowered if you have a coupon. Here they increase the price (because the buyer is using a credit card). In this case the list price $22 950 is increased to 24097.50

OpenStudy (anonymous):

@phi thanks :)